# A Robust Adversarial Ensemble with Causal (Feature Interaction) Interpretations for Image Classification

### v1

@@ Abstract
pattern recognition, image classification, adversarial attacks, generative classifier
@@ 1 Introduction
In recent years, various deep learning-based discriminative models have made significant progress in image classification [1, 2], such as VGG [3], ResNet [4], and ViT [5]. Despite their great success, recent studies have uncovered that deep discriminative classifiers are vulnerable to adversarial examples [6, 7, 8, 9]. These are well-designed inputs with slight perturbations that can cause the model to produce incorrect classification results. This vulnerability partly stems from the opaque nature of the black-box deep neural network (DNN) models and their limited interpretability. Adversarial examples are typically modified slightly so that humans don’t misclassify them, thus posing a non-trivial threat to many DNN-based applications.
Adversarial training has emerged as one of the most effective techniques for improving robustness [7, 10, 11]. The core idea is to include adversarial examples in the training stage. However, training data augmentation doesn’t address the root cause of adversarial vulnerability: the lack of model interpretability. Recent studies show that deep generative classifiers, despite their limited performance on image classification tasks [12, 13, 14], can be more robust to adversarial examples [15]. This adversarial robustness benefits from their ability to model the distribution of each class, essentially modeling how a specific class generates the input data. This approach leads to better model interpretability and allows them to better understand what adversarial inputs might look like. Discriminative classifiers, on the other hand, model decision boundaries between different classes and learn which input features contribute most to distinguishing between various classes. Consequently, they are more vulnerable to adversarial examples designed to create outliers for decision boundaries, thus confusing the classifier. This difference provides the underlying motivation and theoretical support for using generative classifiers to build more adversarially robust networks, as it’s more challenging to shift feature distribution than to create outliers.
@@ 2 Related Work
Adversarial training has emerged as an effective approach to enhance model robustness. Goodfellow et al. [7] and Kurakin et al. [16] demonstrated significant error rate reductions on MNIST and ImageNet datasets, respectively, through adversarial training. However, injecting adversarial examples into the training set can decrease accuracy on clean datasets, and the process is computationally expensive due to the construction of sophisticated adversarial examples [16]. To address these issues, Shafahi et al. [10] introduced a ”free” adversarial training algorithm that eliminates the overhead cost of generating adversarial examples by recycling the gradient information computed when updating model parameters. Recent improvements in adversarial training include using cyclic learning rates to accelerate training [11] and employing diffusion models to augment adversarial training datasets [17]. Nevertheless, it remains unrealistic to involve all types of adversarial examples in the training phase due to the variety of attacks. More importantly, adversarial training doesn’t solve the intrinsic problem of adversarial vulnerability, which lies in the lack of model interpretability.
@@
Recent research has begun to evaluate the adversarial robustness of generative classifiers, which offer improved interpretability. Li et al. [15] found that deep generative classifiers can be robust on MNIST and CIFAR-binary datasets. Zhang et al. [18] further improved deep generative classifier robustness by modeling adversarial perturbation from a causal perspective. However, both studies were unable to obtain satisfactory results on full image datasets, reporting that VAE-based generative classifiers achieved less than 50% clean test accuracy on CIFAR-10. Additionally, they didn’t conduct an interpretability analysis to show the correlations between adversarial robustness and model interpretability. Mackowiak et al. [12] showed that Invertible Neural Networks (INNs) are more interpretable than ResNet-50. However, they demonstrated that INNs do not fully prevent C&W attacks [8] on ImageNet, despite C&W attacks prioritizing small perturbations over attack success rates.
@@
To address these limitations, we develop a generalized deep generative classifier architecture for more complex datasets (full CIFAR-10 and CIFAR-100), using only clean data for training to reduce training time and computational costs. To demonstrate the generalizability of our approach, we conduct preliminary experiments on the Tiny-ImageNet dataset, showcasing the potential of our method to scale to larger datasets for real-world applications. Furthermore, we evaluate robustness against one of the strongest attacks (PGD [9]), which prioritizes attack success rates over perturbation size. Importantly, we explore the correlations between adversarial robustness and model interpretability, showing that by designing causal graphs in generative classifiers, we can achieve better model interpretability, leading to improved adversarial robustness.
@@ 3 Methodology
Our work extends the approach presented in [18] to handle more complex image classification tasks (e.g. full CIFAR-10, CIFAR-100 and Tiny-ImageNet datasets). The top-level generative classifier processes both the features extracted from the bottom-level pre-trained CNN and the original image. This dual-input approach leverages the generative model’s ability to regenerate its inputs, enabling it to learn distributions from both feature representations and original data. Such comprehensive learning strategy enhances the model’s capability to handle both adversarial images and adversarial features.
@@ 3.1 Latent Variable Model
Following Goodfellow et al. [7], we model adversarial perturbations $M$ as a specific type of noise affecting both $X_{1}$ and $X_{2}$. In white-box attacks, these perturbations $M$ are generated based on the labels $Y$, input data $(X_{1},X_{2})$, and network parameters $\theta$. The formation of adversarial inputs $(X_{1},X_{2})$ is then caused by the combination of adversarial perturbations $M$, labels $Y$, and other factors $Z$. For simplicity, we do not consider cases where other factors $Z$ influence labels $Y$. The causal model for input data formation can be expressed as:
@@ 3.2 VAE-based Generative Classifier
Building upon the causal relations established in Section 3.1, we develop our generative classifier using amortized variational inference [19]. The fundamental approach applies Bayes’ theorem to estimate the conditional probability $p(y|x)$, where $y$ is the target label and $x$ is the input. A generative classifier predicts the label $y$ of an input $x$ according to:
@@
[…]
Due to the intractability of the integral of the marginal log-likelihood, arising from intractable true posterior densities $p(z|\cdot)$ and $p(m|\cdot)$ for latent variables, we introduce an approximate distribution $q(z,m;\lambda)$ with variational parameters $\lambda$ to approximate the true posterior [20]. Then the training objective of maximizing log-likelihood function in Eq. (5) is equivalent to minimizing the Kullback-Leibler (KL) divergence [21] between the variational distribution and true posterior $D_{KL}(q(z,m;\lambda)||p(z,m|\cdot))$. However, this divergence $D_{KL}$ is almost impossible to minimize to zero because the variational distribution is usually not capable enough to catch the complexity of the true posterior due to insufficient parameters. To address this challenge, we maximize the Evidence Lower Bound (ELBO), which is equivalent to minimizing the divergence [19]. ELBO is a lower bound on the log marginal probability of the data and can be derived from Eq. (5) using Jensen’s inequality:
[…]
@@
We initialize $p(z)$ as a Gaussian distribution ($\mu=0$, $\sigma=0$) and $p(y)$ as a uniform distribution based on the dataset’s classes (i.e., 0.1 for CIFAR-10 and 0.01 for CIFAR-100). During inference, $m$ is not set to 00 but sampled from $q_{\lambda_{2}}(m|x_{1},x_{2},y_{c})$ and $z$ is sampled from $q_{\lambda_{1}}(z|x_{1},x_{2},y_{c},m)$. The final prediction is:
@@ 4.1 Experimental Setup
Datasets. Generative classifier benchmarking typically focuses on MNIST, SVHN, or CIFAR datasets due to performance constraints [12, 14]. In this paper, we utilize the full CIFAR-10 and CIFAR-100 [22] for comprehensive evaluation of both clean accuracy and adversarial robustness. Additionally, we conduct clean accuracy experiments on Tiny-ImageNet [23] to demonstrate the generalizability of our approach.
Feature extractors. We employ VGG-16 and VGG-19 [3] as our pre-trained feature extractors. Despite not being the most recent models, VGG remains widely used as backbones in state-of-the-art (SOTA) CNN and transformer-based vision models [24, 25] and is capable enough to deal with CIFAR and Tiny-ImageNet data. Rather than using the entire VGG model, we extract features from intermediate layers. We implement a shallow model (containing only convolution layers) and a deep model (containing both convolution and dense layers) for ablation study, with further details discussed in Section 4.4. Our bottom-up architecture design also allows seamless integration of alternative backbones, such as ResNet [4], without requiring modifications to the generative component.
@@
Adversarial attacks. Following the evaluation framework in [15, 18], we target white-box attacks and evaluate our model against Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD). FGSM [7] is selected for its simplicity and representativeness as a classic attack. PGD [9] represents SOTA attack performance and is consistently ranked among the strongest attacks [26, 27]. For both attacks, we evaluate perturbation magnitudes $\epsilon$ ranging from 00 to $0.2$, where $\epsilon=0$ represents clean inputs. The perturbation size is constrained to maintain visual similarity between original and adversarial inputs. We employ the $l_{2}$-norm with $30$ steps for PGD attacks.
Interpretability. We use counterfactual explanations to evaluate interpretability, treating adversarial examples as a specific type of counterfactuals. We evaluate these based on proximity (perturbation size) and speed (attack iterations) [28]. Lower proximity and higher speed indicate higher quality counterfactuals. Models more susceptible to generating high-quality counterfactuals demonstrate lower interpretability, and this indicates greater vulnerability to adversarial attacks.
Additionally, we employ feature attribution analysis through Remove and Retrain (ROAR) [29]. Using integrated gradients [30] for feature attribution, we replace 30% of the most important pixels with the mean of CIFAR data, followed by model retraining and re-evaluation on the modified dataset. Greater accuracy degradation in this case indicates better interpretability, as it shows the model’s reliance on more important features.
@@ 4.2 Evaluation of Adversarial Robustness
We evaluate our model’s robustness against adversarial attacks as shown in Fig. 4(a), 4(d), 5(a), and 5(d). The baseline models include three adversarially robustified generative classifiers (GBZ-CONV9 [15], GBY-CONV9 [15], and Deep CAMA [18]) and one discriminative classifier (DBX-FC1 [15]). Additionally, as our model employs VGG16 as a feature extractor, we include a pure VGG-based discriminative classifier for direct comparison. Results from both shallow and deep models are reported and ablation study is shown in Section 4.4.
Our proposed model demonstrates superior adversarial robustness across different attack types and strengths. For FGSM attacks, the model maintains remarkable robustness on both CIFAR-10 and CIFAR-100 datasets, with accuracy degradation limited to less than $0.1$, effectively mitigating the FGSM attack’s impact. Against the more challenging PGD attack, our model achieves impressive results on CIFAR-10, maintaining over $0.8$ accuracy at $\epsilon=0.05$ and $0.5$ accuracy at $\epsilon=0.1$, representing a significant robustness improvement. While the PGD attack results on CIFAR-100 show reduced effectiveness compared to CIFAR-10, our model still outperforms all baseline models. The generative classifiers GBZ-CONV9 and GBY-CONV9 also show robustness on both attacks and datasets, but the improvement is limited compared to our approach. Notably, Deep CAMA fails to generalize effectively to CIFAR-10 and CIFAR-100, achieving clean accuracy below $0.5$ on both datasets, consistent with the limitations reported in their paper [18]. The discriminative baselines (VGG16 and DBX-FC1) exhibit the worst robustness across both attacks and datasets. Their performance degrades significantly even with minimal perturbations ($\epsilon<0.05$), with accuracy falling below $0.2$. These results highlight the advantages of our hybrid approach over purely discriminative or generative classifiers.
Additionally, our model maintains high clean dataset accuracy ($\epsilon=0$) without the typical degradation observed in adversarial training [16]. We attribute this to the enhanced interpretability (showed in Section 4.3), which appears to contribute to overall model performance. Traditional adversarial training typically requires 3-30 times longer training periods compared to non-robust equivalents [10]. While our study doesn’t include quantitative training time comparisons with adversarially trained baselines, our approach provides inherent efficiency advantages as it achieves robustness without dataset augmentation or enlargement through injecting adversarial examples.
@@ 4.3 Evaluation of Model Interpretability
VGG16 [3] 0.4828
DBX-FC1 [15] 0.4962
Deep CAMA [18] 0.2462
GBZ-CONV9 [15] 0.4278
GBY-CONV9 [15] 0.5571
@@
Tiny- ImageNet Shallow (14 CONV) 0.4734

### v2

# Causal Interpretability for Adversarial Robustness: A Hybrid Generative Classification Approach

@@ Abstract
pattern recognition, image classification, adversarial attack, generative classifier
@@ 1 Introduction
In recent years, various deep learning-based discriminative models have made significant progress in image classification [mauricio2023comparing, wang2021comparative], such as VGG [simonyan2014very], ResNet [he2016deep], and ViT [dosovitskiy2020image]. Despite their great success, recent studies have uncovered that deep discriminative classifiers are vulnerable to adversarial examples [szegedy2013intriguing, goodfellow2014explaining, carlini2017towards, madry2017towards]. These are well-designed inputs with slight perturbations that can cause the model to produce incorrect classification results. This vulnerability partly stems from the opaque nature of the black-box deep neural network (DNN) models and their limited interpretability. Adversarial examples are typically modified slightly so that humans don’t misclassify them, thus posing a non-trivial threat to many DNN-based applications.
Adversarial training has emerged as one of the most effective techniques for improving robustness [goodfellow2014explaining, shafahi2019adversarial, wong2020fast]. The core idea is to include adversarial examples in the training stage. However, training data augmentation doesn’t address the root cause of adversarial vulnerability: the lack of model interpretability. Recent studies show that deep generative classifiers, despite their limited performance on image classification tasks [mackowiak2021generative, fetaya2019understanding, zheng2023revisiting], can be more robust to adversarial examples [li2019generative]. This adversarial robustness benefits from their ability to model the distribution of each class, essentially modeling how a specific class generates the input data. This approach leads to better model interpretability and allows them to better understand what adversarial inputs might look like. Discriminative classifiers, on the other hand, model decision boundaries between different classes and learn which input features contribute most to distinguishing between various classes. Consequently, they are more vulnerable to adversarial examples designed to create outliers for decision boundaries, thus confusing the classifier. This difference provides the underlying motivation and theoretical support for using generative classifiers to build more adversarially robust networks, as it’s more challenging to shift feature distribution than to create outliers.
@@ 2 Related Work
Adversarial training has emerged as an effective approach to enhance model robustness. Goodfellow et al. [goodfellow2014explaining] and Kurakin et al. [kurakin2016adversarial] demonstrated significant error rate reductions on MNIST and ImageNet datasets, respectively, through adversarial training. However, injecting adversarial examples into the training set can decrease accuracy on clean datasets, and the process is computationally expensive due to the construction of sophisticated adversarial examples [kurakin2016adversarial]. To address these issues, Shafahi et al. [shafahi2019adversarial] introduced a ”free” adversarial training algorithm that eliminates the overhead cost of generating adversarial examples by recycling the gradient information computed when updating model parameters. Recent improvements in adversarial training include using cyclic learning rates to accelerate training [wong2020fast] and employing diffusion models to augment adversarial training datasets [wang2023better]. Nevertheless, it remains unrealistic to involve all types of adversarial examples in the training phase due to the variety of attacks. More importantly, adversarial training doesn’t solve the intrinsic problem of adversarial vulnerability, which lies in the lack of model interpretability.
@@
Recent research has begun to evaluate the adversarial robustness of generative classifiers, which offer improved interpretability. Li et al. [li2019generative] found that deep generative classifiers can be robust on MNIST and CIFAR-binary datasets. Zhang et al. [zhang2020causal] further improved deep generative classifier robustness by modeling adversarial perturbation from a causal perspective. However, both studies were unable to obtain satisfactory results on full image datasets, reporting that VAE-based generative classifiers achieved less than 50% clean test accuracy on CIFAR-10. Additionally, they didn’t conduct an interpretability analysis to show the correlations between adversarial robustness and model interpretability. Mackowiak et al. [mackowiak2021generative] showed that Invertible Neural Networks (INNs) are more interpretable than ResNet-50. However, they demonstrated that INNs do not fully prevent C&W attacks [carlini2017towards] on ImageNet, despite C&W attacks prioritizing small perturbations over attack success rates.
@@
To address these limitations, we develop a generalized deep generative classifier architecture for more complex datasets (full CIFAR-10 and CIFAR-100), using only clean data for training to reduce training time and computational costs. To demonstrate the generalizability of our approach, we conduct preliminary experiments on the Tiny-ImageNet dataset, showcasing the potential of our method to scale to larger datasets for real-world applications. Furthermore, we evaluate robustness against one of the strongest attacks (PGD [madry2017towards]), which prioritizes attack success rates over perturbation size. Importantly, we explore the correlations between adversarial robustness and model interpretability, showing that by designing causal graphs in generative classifiers, we can achieve better model interpretability, leading to improved adversarial robustness.
@@ 3 Methodology
Our work extends the approach presented in [zhang2020causal] to handle more complex image classification tasks (e.g. full CIFAR-10, CIFAR-100 and Tiny-ImageNet datasets). The top-level generative classifier processes both the features extracted from the bottom-level pre-trained CNN and the original image. This dual-input approach leverages the generative model’s ability to regenerate its inputs, enabling it to learn distributions from both feature representations and original data. Such comprehensive learning strategy enhances the model’s capability to handle both adversarial images and adversarial features.
@@ 3.1 Latent Variable Model
Following Goodfellow et al. [goodfellow2014explaining], we model adversarial perturbations $M$ as a specific type of noise affecting both $X_{1}$ and $X_{2}$. In white-box attacks, these perturbations $M$ are generated based on the labels $Y$, input data $(X_{1},X_{2})$, and network parameters $\theta$. The formation of adversarial inputs $(X_{1},X_{2})$ is then caused by the combination of adversarial perturbations $M$, labels $Y$, and other factors $Z$. For simplicity, we do not consider cases where other factors $Z$ influence labels $Y$. The causal model for input data formation can be expressed as:
@@ 3.2 VAE-based Generative Classifier
Building upon the causal relations established in Section 3.1, we develop our generative classifier using amortized variational inference [zhang2018advances]. The fundamental approach applies Bayes’ theorem to estimate the conditional probability $p(y|x)$, where $y$ is the target label and $x$ is the input. A generative classifier predicts the label $y$ of an input $x$ according to:
@@
$\begin{split}&p(y|x)=softmax_{c=1}^{C}\\ &[\log\int p(x_{1},x_{2}|y,z,m)p(m)p(z)p(y)\,dm\,dz]\end{split}$ (5)
Due to the intractability of the integral of the marginal log-likelihood, arising from intractable true posterior densities $p(z|\cdot)$ and $p(m|\cdot)$ for latent variables, we introduce an approximate distribution $q(z,m;\lambda)$ with variational parameters $\lambda$ to approximate the true posterior [kingma2013auto]. Then the training objective of maximizing log-likelihood function in Eq. (5) is equivalent to minimizing the Kullback-Leibler (KL) divergence [kullback1951information] between the variational distribution and true posterior $D_{KL}(q(z,m;\lambda)||p(z,m|\cdot))$. However, this divergence $D_{KL}$ is almost impossible to minimize to zero because the variational distribution is usually not capable enough to catch the complexity of the true posterior due to insufficient parameters. To address this challenge, we maximize the Evidence Lower Bound (ELBO), which is equivalent to minimizing the divergence [zhang2018advances]. ELBO is a lower bound on the log marginal probability of the data and can be derived from Eq. (5) using Jensen’s inequality:
$\begin{split}&\log p(x,y)=\log\int p(x_{1},x_{2}|y,z,m)p(m)p(z)p(y)\,dm\,dz\\ &=\log\int p(x_{1},x_{2}|y,z,m)p(m)p(z)p(y)\frac{q(z,m;\lambda)}{q(z,m;\lambda)}\,dm\,dz\\ &=\log\mathbb{E}_{q(z,m;\lambda)}\left[\frac{p(x_{1},x_{2}|y,z,m)p(m)p(z)p(y)}{q(z,m;\lambda)}\right]\\ &\geq\mathbb{E}_{q(z,m;\lambda)}\left[\log\frac{p(x_{1},x_{2}|y,z,m)p(m)p(z)p(y)}{q(z,m;\lambda)}\right]\end{split}$ (6)
@@
We initialize $p(z)$ as a Gaussian distribution ($\mu=0$, $\sigma=0$) and $p(y)$ as a uniform distribution based on the dataset’s classes (i.e., 0.1 for CIFAR-10 and 0.01 for CIFAR-100). During inference, $m$ is not set to 0 but sampled from $q_{\lambda_{2}}(m|x_{1},x_{2},y_{c})$ and $z$ is sampled from $q_{\lambda_{1}}(z|x_{1},x_{2},y_{c},m)$. The final prediction is:
@@ 4.1 Experimental Setup
Datasets. Generative classifier benchmarking typically focuses on MNIST, SVHN, or CIFAR datasets due to performance constraints [mackowiak2021generative, zheng2023revisiting]. In this paper, we utilize the full CIFAR-10 and CIFAR-100 [krizhevsky2009learning] for comprehensive evaluation of both clean accuracy and adversarial robustness. Additionally, we conduct clean accuracy experiments on Tiny-ImageNet [le2015tiny] to demonstrate the generalizability of our approach.
Feature extractors. We employ VGG-16 and VGG-19 [simonyan2014very] as our pre-trained feature extractors. Despite not being the most recent models, VGG remains widely used as backbones in state-of-the-art (SOTA) CNN and transformer-based vision models [hu2023vgg, amjoud2023object] and is capable enough to deal with CIFAR and Tiny-ImageNet data. Rather than using the entire VGG model, we extract features from intermediate layers. We implement a shallow model (containing only convolution layers) and a deep model (containing both convolution and dense layers) for ablation study, with further details discussed in Section 4.4. Our bottom-up architecture design also allows seamless integration of alternative backbones, such as ResNet [he2016deep], without requiring modifications to the generative component.
@@
Adversarial attacks. Following the evaluation framework in [li2019generative, zhang2020causal], we target white-box attacks and evaluate our model against Fast Gradient Sign Method (FGSM) and Projected Gradient Descent (PGD). FGSM [goodfellow2014explaining] is selected for its simplicity and representativeness as a classic attack. PGD [madry2017towards] represents SOTA attack performance and is consistently ranked among the strongest attacks [xue2024diffusion, wang2020adversarial]. For both attacks, we evaluate perturbation magnitudes $\epsilon$ ranging from 0 to $0.2$, where $\epsilon=0$ represents clean inputs. The perturbation size is constrained to maintain visual similarity between original and adversarial inputs. We employ the $l_{2}$-norm with $30$ steps for PGD attacks.
Interpretability. We use counterfactual explanations to evaluate interpretability, treating adversarial examples as a specific type of counterfactuals. We evaluate these based on proximity (perturbation size) and speed (attack iterations) [moraffah2020causal]. Lower proximity and higher speed indicate higher quality counterfactuals. Models more susceptible to generating high-quality counterfactuals demonstrate lower interpretability, and this indicates greater vulnerability to adversarial attacks.
Additionally, we employ feature attribution analysis through Remove and Retrain (ROAR) [hooker2019benchmark]. Using integrated gradients [sundararajan2017axiomatic] for feature attribution, we replace 30% of the most important pixels with the mean of CIFAR data, followed by model retraining and re-evaluation on the modified dataset. Greater accuracy degradation in this case indicates better interpretability, as it shows the model’s reliance on more important features.
@@ 4.2 Evaluation of Adversarial Robustness
We evaluate our model’s robustness against adversarial attacks as shown in Fig. 4(a), 4(d), 5(a), and 5(d). The baseline models include three adversarially robustified generative classifiers (GBZ-CONV9 [li2019generative], GBY-CONV9 [li2019generative], and Deep CAMA [zhang2020causal]) and one discriminative classifier (DBX-FC1 [li2019generative]). Additionally, as our model employs VGG16 as a feature extractor, we include a pure VGG-based discriminative classifier for direct comparison. Results from both shallow and deep models are reported and ablation study is shown in Section 4.4.
Our proposed model demonstrates superior adversarial robustness across different attack types and strengths. For FGSM attacks, the model maintains remarkable robustness on both CIFAR-10 and CIFAR-100 datasets, with accuracy degradation limited to less than $0.1$, effectively mitigating the FGSM attack’s impact. Against the more challenging PGD attack, our model achieves impressive results on CIFAR-10, maintaining over $0.8$ accuracy at $\epsilon=0.05$ and $0.5$ accuracy at $\epsilon=0.1$, representing a significant robustness improvement. While the PGD attack results on CIFAR-100 show reduced effectiveness compared to CIFAR-10, our model still outperforms all baseline models. The generative classifiers GBZ-CONV9 and GBY-CONV9 also show robustness on both attacks and datasets, but the improvement is limited compared to our approach. Notably, Deep CAMA fails to generalize effectively to CIFAR-10 and CIFAR-100, achieving clean accuracy below $0.5$ on both datasets, consistent with the limitations reported in their paper [zhang2020causal]. The discriminative baselines (VGG16 and DBX-FC1) exhibit the worst robustness across both attacks and datasets. Their performance degrades significantly even with minimal perturbations ($\epsilon<0.05$), with accuracy falling below $0.2$. These results highlight the advantages of our hybrid approach over purely discriminative or generative classifiers.
Additionally, our model maintains high clean dataset accuracy ($\epsilon=0$) without the typical degradation observed in adversarial training [kurakin2016adversarial]. We attribute this to the enhanced interpretability (showed in Section 4.3), which appears to contribute to overall model performance. Traditional adversarial training typically requires 3-30 times longer training periods compared to non-robust equivalents [shafahi2019adversarial]. While our study doesn’t include quantitative training time comparisons with adversarially trained baselines, our approach provides inherent efficiency advantages as it achieves robustness without dataset augmentation or enlargement through injecting adversarial examples.
@@ 4.3 Evaluation of Model Interpretability
VGG16 [simonyan2014very] 0.4828
DBX-FC1 [li2019generative] 0.4962
Deep CAMA [zhang2020causal] 0.2462
GBZ-CONV9 [li2019generative] 0.4278
GBY-CONV9 [li2019generative] 0.5571
@@
Tiny-
ImageNet
Shallow (14 CONV) 0.4734